OpenStudy (anonymous):

i = $125 r = 6% t = 1 What is p? $56.25 $2,027.08 $2,083.33

OpenStudy (anonymous):

okay now with this one i did what i did last time and didnt get the answer?

OpenStudy (anonymous):

Maybe you did it wrong the last time as well.

OpenStudy (anonymous):

What is the equation that you use to solve for p?

OpenStudy (anonymous):

well i multiplied it

OpenStudy (anonymous):

Obviously that's wrong. What equation does your book give you?

OpenStudy (anonymous):

125times6%times1=7.5

OpenStudy (anonymous):

Ok then the answer is 7.5

OpenStudy (anonymous):

its an online lesson no book

OpenStudy (anonymous):

you work backwards here

OpenStudy (anonymous):

You can't work any way because we have no equation

OpenStudy (anonymous):

you already have i so you take your % and multiply each one of the given answers to get 125

OpenStudy (anonymous):

do you understand what im saying?

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

Can someone give me the equation?

OpenStudy (anonymous):

i = r * t * p

OpenStudy (anonymous):

Ok we want to find P so we isolate P in the equation i = r * t * p i /r*t= p ----We divide r and t to both sides.

OpenStudy (anonymous):

Now we have the equation to solve for p (i)/(r*t) = p

OpenStudy (anonymous):

You are given i r t plug them in and solve it